Write an equation of the line that is parallel to 2x + 4y = 6 and passes through the point (6, 4).

Answer by mananth(16946)

Parallel to 2x+4y=6 and passing through(6,4)
2 x 4 y = 6
Find the slope of this line
make y the subject
4 y = -2 x + 6
Divide by 4
y = - 1/2 x + 3/2
Compare this equation with y=mx+b
slope m = - 1/2
The slope of a line parallel to the above line will be the same
The slope of the required line will be - 1/2
m= - 1/2 ,point ( 6 , 4 )
Find b by plugging the values of m & the point in
y=mx+b
4 = -3 + b
b= 7
m= - 1/2
Plug value of the slope and b in y = mx +b
The required equation is y = - 1/2 x + 7

m.ananth@hotmail.ca